Mir Status Report 12 January 1998

NASA MIR Program / Mir 24 / NASA 6 Status
Date: Monday, January 12, 1998 // MD 160/109

Mir Station System/Subsystem Activities
The crew performed separation on the Base Block cooling loops KOB-1 and KOB-2 this afternoon after flight controllers detected air in those loops. All other systems continue to operate nominally as previously reported.

The pressure in the primary airlock (ShSO) has decreased to below 570 mmHg. Russian ground teams are continuing to investigate the situation to determine the correct course of action following the upcoming EVA.

Plans remain on schedule for EVA #7 this week. EVA times are currently shown as:

Crew wake DMT 14 Jan 16:45
Hatch Open DMT 14 Jan 23:40
Hatch Close DMT 15 Jan 02:15
Crew sleep DMT 15 Jan 07:45

PAYLOAD ACTIVITIES
Space Suit Prep
SMP:ENV:I:48 Activate SSAS
SMP:ENV:I:49 Grab Air Sample
SMP:MO:I:56 FMK - Activate Personal/Area Monitors
SMP:MO:I:56 FMK - Deactivate Personal Monitor
FB:HSR:I:23 CHAPAT Daily Check

Data Management
On Board -All MIPS files received on the 11th were corrupted. MOST believes there was a problem with the ground site. Dave has been asked to resend all of those files.

SMP/Medical -- Dave performed Air Quality sampling today. This included the activation of the Solid Sorbent Air Samplers (SSAS) and Formaldehyde Monitoring. Grab Samples were completed in five locations. Dave reported that one GSC did not have a vacuum so he used a spare. This leaves two GCS's onboard for the pre- and post-docking time period.

Environmental Parameters (last 24 hours):
Parameter Min. Max. Nominal Range
P (mmHg) BB 668 671 660 - 850
ppO2 (mmHg) 151 153 140 - 200
ppCO2 (mmHg) 1.6 2.2 up to 8.0
ppH20 (mmHg) 11.7 12.7 8 - 14
Humidity (%) 38 43 30 - 75
Temp (xC) BB 28.3 29.6 18 - 28

Communications Summary
13:35 - 13:38 EVA, COSS

Planning
MIR 24/NASA 6 MD 161/110
Suited Dry Run
SMP:ENV:I:50 Deactivate SSAS
SMP:MO:I:56 FMK - Deactivate Area Monitor
SMP:ENV:I:70 CPA Measurement
OPS:FAC:I:11 MIPS D/L and Prep for Downlink

MIR 24/NASA 6 MD 162/111
FB:HSR:I:23 CHAPAT Daily Check
OPS:FAC:I:11 MIPS D/L and Prep for Downlink
HLS:BP:I:10 Interactions Questionnaire Mir Crew
HLS:BP:I:13 Interactions Questionnaire - NASA Crew
ISS:ENV:I:64 SPSR Assembly
ISS:ENV:I:65 SPSR Checkout CREAM PDP Relocation

  • submit to reddit


Loading



MILITARY SPACE USA, 11TH JUNE TO 12TH JUNE 2019, LOS ANGELES, USA
49th International Conference on Environmental Systems - ICES 2019




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on January 12, 1998 9:30 PM.

ALERT! Paradigm Shift Ahead - Has Dan Goldin sent Human Space Flight off in a new direction? was the previous entry in this blog.

House Science Committee Hearing Summary is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.